A triangle lakefront lot has a perimeter of 600 feet. One side is 50 feet shorter than 2 times the shortest side, while the third side is 150 feet longer than the shortest side.
Find the lengths of all three sides?

Answer by LinnW(1048)

let s = short side
m = medium side
l = long side
m = 2s - 50
l = 150 + s
perimeter = s + m + l
600 = s + m + l
substitute (2s - 50) for m, and (150 + s) for l
600 = s + (2s - 50) + (150 + s)
600 = 4s + 100
add -100 to each side
500 = 4s
divide each side by 4
125 = s
m = 2(125) - 50 = 250 - 50 = 200
l = 150 + s = 150 + 125 = 275
So the side lengths are 125,200 and 275 .